# CoSy: Evaluating Textual Explanations of Neurons

### v1

@@ A.1.1 NetDissect
Network Dissection (NetDissect) [Bau_2017_CVPR] is a method designed to explain individual neurons of DNNs, particularly convolutional neural networks (CNNs) within the domain of CV. This approach systematically analyzes the network’s learned concepts by aligning individual neurons with given semantic concepts. To perform this analysis, annotated datasets with segmentation masks are required, where these masks label each pixel in an image with its corresponding object or attribute identity. The Broadly and Densely Labeled Dataset (Broden) [Bau_2017_CVPR] combines a set of densely labeled image datasets that represent both low-level concepts, such as colors, and higher-level concepts, such as objects. It provides a comprehensive set of ground truth examples for a broad range of visual concepts such as objects, scenes, object parts, textures, and materials in a variety of contexts.
A concept $s\in\mathcal{C}\subset\mathcal{S}$ is defined as a visual concept in NetDissect and is provided by the pixel-level annotated Broden dataset. Given a CNN and the Broden dataset as input, NetDissect explains a neural representation $f_{i}$ by searching for the highest similarity between concept image segmentation masks and neuron activation masks. Concept image segmentation masks are provided by the Broden dataset $B_{s}({\boldsymbol{x}})\in\{0,1\}^{H\times W}$, where a value of 1 signifies the pixel-level presence of $s$, and 0 denotes its absence. Neuron activation masks are obtained by thresholding the continuous neuron activations of $f_{i}$ into binary masks $A({\boldsymbol{x}})\in\{0,1\}^{H\times W}$. Then the similarity $\delta_{\text{IoU}}$ between image segmentation masks and binary neuron masks can be evaluated using the Intersection over Union score (IoU) for an individual neuron within a layer:
@@ A.1.2 CompExp
To overcome the limitation of explaining neurons with only a single concept, the Compositional Explanations of Neurons (CompExp) method was later introduced [muCompositionalExplanationsNeurons2021], enabling the labeling of neurons with compositional concepts. The method obtains its explanations by merging individual concepts into logical formulas using composition operators AND, OR, and NOT. The formula length $L\in\mathbb{N}$ is defined beforehand. The initial stage of explanation generation is similar to NetDissect, a set of images is taken as input, and convolutional neuron activations are converted into binary masks. The explanations are constructed through a beam search algorithm [cormen2022introduction], beginning with individual concepts and gradually building them into more complex logical formulas. Throughout the beam search stages, the existing formulas in the beam are combined with new concepts. These new formulas are measured by the IoU. The maximization of the IoU score is desired to get a high explanation quality.
@@ A.1.3 MILAN
MILAN [hernandez2022natural] is a method that aims to describe neural representations within a DNN through open-ended natural language descriptions. First, a dataset of fine-grained human descriptions of image regions (Milannotations) is collected. These descriptions can be defined as concepts that are open-ended natural language descriptions, where $s\in\mathcal{N}\subset\mathcal{S}$. Given a DNN and input images $\boldsymbol{x}\in\boldsymbol{X}$, neuron masks $M({\boldsymbol{x}})\in\mathbb{R}^{H\times W\times C}$ are collected of highly activated image regions for $f_{i}$.
Two distributions are then derived: the probability $p(s|M({\boldsymbol{x}}))$ that a human would describe an image region with $s$, and the probability $p(s)$ that a human would use the description $s$ for any neuron. The probability $p(s|M({\boldsymbol{x}}))$ is approximated with the Show-Attend-Tell [xu2015show] image-to-text model trained on the Milannotations dataset. Additionally, $p(s)$ is approximated with a two-layer LSTM language model [hochreiter1997long] trained on the Milannotations dataset.
@@ A.1.4 FALCON
The FALCON [pmlr-v202-kalibhat23a] explainability method has a similar approach to MILAN. Initially, it gathers the most highly activating images corresponding to a neural representation. GradCam [selvaraju2017grad] is subsequently applied to identify highlighted features in these images, which are then cropped to focus on these regions. These cropped images, along with large captioning dataset LAION-400m [Schuhmann2021Laion] with concepts $s\in\mathcal{N}\subset\mathcal{S}$, are input to CLIP (Contrastive Language-Image Pre-training) [radford2021learning], which computes the image-text similarity between the text embeddings of captions and the input cropped images. The top 5 captions are then extracted. Conversely, the least activating images are collected, and concepts are extracted and removed from the top-scoring concepts, ultimately yielding the explanation of the neural representation.
@@
These parameter settings significantly restrict the number of explainable neurons, resulting to fewer than 50 explainable neurons. This constraint prevents the necessary randomization for comparison with other methods. To address this, we set $\alpha=0$ and $\gamma=0$. However, for FALCON-original, we retain the original settings of $\alpha$ and $\gamma$ and calculate $\Psi$ across all “explainable neurons.” In our experiments on ResNet18, FALCON can only be applied to layers 2 to 4.
@@ A.1.5 CLIP-Dissect
CLIP-Dissect [oikarinen2022clip] is an explanation method that describes neurons in vision DNNs with open-ended concepts, eliminating the need for labeled data or human examples. This method integrates CLIP [radford2021learning], which efficiently learns deep visual representations from natural language supervision. It utilizes both the image encoder and text encoder components of a CLIP model to compute the text embedding for each concept $s\in\mathcal{C}\subset\mathcal{S}$ from a concept dataset and the image embeddings for the probing images in the dataset, subsequently calculating a concept-activation matrix.
The activations of a target neuron $f_{i}$ are then computed across all images in the probing dataset $\boldsymbol{X}$. However, as this process is designed for scalar neural representations, these activations are summarized by a function that calculates the mean of the activation map over spatial dimensions. The concept corresponding to the target neuron is determined by identifying the most similar concept $s$ based on its activation vector. The most highly activated images are denoted as $\boldsymbol{X}_{s}\subset\boldsymbol{X}$.
@@ A.1.6 INVERT
Labeling Neural Representations with Inverse Recognition (INVERT) [bykov2023labeling] shares the capability of constructing complex explanations like CompExp [muCompositionalExplanationsNeurons2021] but with the added advantage of not relying on segmentation masks and only needing labeled data. The method obtains its explanations by merging individual concepts into logical formulas using composition operators AND, OR, and NOT. It also exhibits greater versatility in handling various neuron types and is computationally less demanding compared to previous methods such as NetDissect [Bau_2017_CVPR] and CompExp [muCompositionalExplanationsNeurons2021]. Additionally, INVERT introduces a transparent metric for assessing the alignment between representations and their associated explanations. The non-parametric Area Under the Receiver Operating Characteristic (AUC) measure evaluates the relationship between representations and concepts based on the representation’s ability to distinguish the presence from the absence of a concept, with statistical significance. The probing dataset with the concept present is labeled as $\boldsymbol{X}_{1}$, while the dataset without the concept is labeled as $\boldsymbol{X}_{0}$.
The goal of INVERT is to identify the concept $s\in\mathcal{L}\subset{S}$ that maximizes $\delta_{\text{AUC}}$ with the neural representation $f_{i}$. Here, $s$ can be a combination of concepts. The optimization process resembles that of CompExp, employing beam search [cormen2022introduction] to find the optimal compositional concept. The top-performing concepts are iteratively selected until the predefined compositional length $L\in\mathbb{N}$ is reached.
@@
The objective of INVERT is to maximize the similarity $\delta_{\text{AUC}}$ between a concept $s$ and the neural representation $f_{i}$, which can be described as:
@@ A.2 Schematic Illustration of CoSy Implementation Details
In the example shown in Figure 1, we used the default settings of the explanation methods to generate explanations for neuron 80 in the avgpool layer of ResNet18. For CLIP-Dissect, we used the 20,000 most common English words as the concept dataset and the ImageNet validation dataset [russakovsky2015imagenet] as the probing dataset . We employed Stable Diffusion XL 1.0-base (SDXL) [podell2023sdxl] as the text-to-image model, using the prompt “realistic photo of a close up of [concept]” to generate concept images, with [concept] being replaced by the textual explanation from the methods. We generated 50 images per concept for 50 randomly chosen neurons from the avgpool layer of ResNet18. For evaluation, we also used the ImageNet validation dataset as the control dataset.
A.3 Compute Resources
A.4 Concept Broadness
While CoSy focuses on measuring the explanation quality, another open question is how broad or abstract are the concepts provided as textual explanations. This question of how specific or general an individual neuron is described by the explanation, might be relevant to different XAI applications. For example, research fields where the user aims to deploy the same network for multiple tasks with varying image domains. In this case, describing a neuron’s more general concept such as “a round object” might be more informative than a more (domain-)specific concept such as “a tennis ball” for the network assessment. In an effort to provide insight on the broadness of concepts, we assessed whether the similarity between images generated based on the same concept changes for more general to more specific concepts.
Figure 7: The figure demonstrates the independence of the concept broadness measured by the number of hypernyms as defined in WordNet [miller1995wordnet] to the inter-image similarity of corresponding generated images.
In our experiment, we define the broadness of a concept based on the number of hypernyms in the WordNet hierarchy [miller1995wordnet]. The more specific a concept the larger the number of hypernyms. We choose two ImageNet classes (“ladybug”, “pug”) and generate $50$ images for each concept as well as each hypernym of both concepts (with the most general concept being “entity”.) Then, we measure the cosine similarity of all images generated based on the same concept. The box plot of the cosine similarity across both concepts and all hypernyms, in Figure 7 indicates that we do not find a correlation. Thus, we hypothesize that the chosen temperature of the diffusion model has a stronger effect on image similarity than the broadness of the prompt used for image generation.
@@ A.5 Intraclass Image Similarity
Cosine similarity (CS) and “Euclidean distance” (ED) are commonly used metrics for measuring image similarity because they capture different aspects of similarity and complement each other. We compute the average CS and ED for each class and determine the overall class average. Table 5 provides a detailed overview of the results quantifying the similarity within synthetic images using CS and ED. When evaluating these results, it is important to note that high scores do not necessarily indicate optimal outcomes, as they suggest nearly identical images, which may lack intraclass distance. Conversely, very low scores imply significant differences among images, which might not capture the essence of the concept adequately. Ideally, we aim for somewhat similar yet slightly varied images representing the same class. The results show that the Stable Cascade (SC) model consistently achieves higher scores across all prompts compared to the Stable Diffusion XL 1.0-base (SDXL) model. Notably, it obtains the highest score for the two most elaborate prompts (4, 5). This indicates that the SC model tends to offer less intraclass distance in visually representing concepts.
Table 5: Intraclass Image Similarity. This table illustrates the impact of varying parameters within our CoSy framework. We evaluate 5 different prompts as input to 2 different text-to-image models. A random selection of 10 ImageNet classes is made, with each class name used as input to the prompt, denoted as [concept], resulting in 50 images generated per prompt using a text-to-image model. We compute the average intraclass similarity across all classes. Higher CS and lower ED values indicate greater similarity between the images. In intraclass image similarity, neither excessively high nor excessively low scores are desirable.
Prompt Text-to-image CS ($\uparrow$) ED ($\downarrow$)
1. “a [concept] SDXL 0.83$\pm$0.07 5.85$\pm$1.41
SC 0.92$\pm$0.03 4.03$\pm$1.00
2. “a painting of [concept] SDXL 0.87$\pm$0.05 4.94$\pm$1.13
SC 0.92$\pm$0.03 3.95$\pm$0.88
3. “photo of [concept] SDXL 0.81$\pm$0.07 6.13$\pm$1.36
SC 0.90$\pm$0.04 4.46$\pm$1.05
4. “realistic photo of [concept] SDXL 0.86$\pm$0.06 5.41$\pm$1.34
SC 0.93$\pm$0.03 3.79$\pm$0.85
5. “realistic photo of a close up of [concept] SDXL 0.88$\pm$0.05 5.09$\pm$1.29
SC 0.93$\pm$0.03 3.95$\pm$0.92

### v2

@@ A.1.1 NetDissect
Network Dissection (NetDissect) [11] is a method designed to explain individual neurons of DNNs, particularly convolutional neural networks (CNNs) within the domain of CV. This approach systematically analyzes the network’s learned concepts by aligning individual neurons with given semantic concepts. To perform this analysis, annotated datasets with segmentation masks are required, where these masks label each pixel in an image with its corresponding object or attribute identity. The Broadly and Densely Labeled Dataset (Broden) [11] combines a set of densely labeled image datasets that represent both low-level concepts, such as colors, and higher-level concepts, such as objects. It provides a comprehensive set of ground truth examples for a broad range of visual concepts such as objects, scenes, object parts, textures, and materials in a variety of contexts.
A concept $s\in\mathcal{C}\subset\mathcal{S}$ is defined as a visual concept in NetDissect and is provided by the pixel-level annotated Broden dataset. Given a CNN and the Broden dataset as input, NetDissect explains a neuron $f_{i}$ by searching for the highest similarity between concept image segmentation masks and neuron activation masks. Concept image segmentation masks are provided by the Broden dataset $B_{s}({\boldsymbol{x}})\in\{0,1\}^{H\times W}$, where a value of 1 signifies the pixel-level presence of $s$, and 0 denotes its absence. Neuron activation masks are obtained by thresholding the continuous neuron activations of $f_{i}$ into binary masks $A({\boldsymbol{x}})\in\{0,1\}^{H\times W}$. Then the similarity $\delta_{\text{IoU}}$ between image segmentation masks and binary neuron masks can be evaluated using the Intersection over Union score (IoU) for an individual neuron within a layer:
@@ A.1.2 CompExp
To overcome the limitation of explaining neurons with only a single concept, the Compositional Explanations of Neurons (CompExp) method was later introduced [12], enabling the labeling of neurons with compositional concepts. The method obtains its explanations by merging individual concepts into logical formulas using composition operators AND, OR, and NOT. The formula length $L\in\mathbb{N}$ is defined beforehand. The initial stage of explanation generation is similar to NetDissect, a set of images is taken as input, and convolutional neuron activations are converted into binary masks. The explanations are constructed through a beam search algorithm [49], beginning with individual concepts and gradually building them into more complex logical formulas. Throughout the beam search stages, the existing formulas in the beam are combined with new concepts. These new formulas are measured by the IoU. The maximization of the IoU score is desired to get a high explanation quality.
@@ A.1.3 MILAN
MILAN [13] is a method that aims to describe neurons within a DNN through open-ended natural language descriptions. First, a dataset of fine-grained human descriptions of image regions (Milannotations) is collected. These descriptions can be defined as concepts that are open-ended natural language descriptions, where $s\in\mathcal{N}\subset\mathcal{S}$. Given a DNN and input images $\boldsymbol{x}\in\boldsymbol{X}$, neuron masks $M({\boldsymbol{x}})\in\mathbb{R}^{H\times W\times C}$ are collected of highly activated image regions for $f_{i}$.
Two distributions are then derived: the probability $p(s|M({\boldsymbol{x}}))$ that a human would describe an image region with $s$, and the probability $p(s)$ that a human would use the description $s$ for any neuron. The probability $p(s|M({\boldsymbol{x}}))$ is approximated with the Show-Attend-Tell [48] image-to-text model trained on the Milannotations dataset. Additionally, $p(s)$ is approximated with a two-layer LSTM language model [50] trained on the Milannotations dataset.
@@ A.1.4 FALCON
The FALCON [14] explainability method has a similar approach to MILAN. Initially, it gathers the most highly activating images corresponding to a neuron. GradCam [5] is subsequently applied to identify highlighted features in these images, which are then cropped to focus on these regions. These cropped images, along with large captioning dataset LAION-400m [51] with concepts $s\in\mathcal{N}\subset\mathcal{S}$, are input to CLIP (Contrastive Language-Image Pre-training) [27], which computes the image-text similarity between the text embeddings of captions and the input cropped images. The top 5 captions are then extracted. Conversely, the least activating images are collected, and concepts are extracted and removed from the top-scoring concepts, ultimately yielding the explanation of the neuron.
@@
These parameter settings significantly restrict the number of explainable neurons, resulting to fewer than 50 explainable neurons. This constraint prevents the necessary randomization for comparison with other methods. To address this, we set $\alpha=0$ and $\gamma=0$. However, for the original FALCON implementation, we retain the original settings of $\alpha$ and $\gamma$ and calculate $\Psi$ across all “explainable neurons”. In our experiments on ResNet18, FALCON can only be applied to layers 2 to 4.
@@ A.1.5 CLIP-Dissect
CLIP-Dissect [52] is an explanation method that describes neurons in vision DNNs with open-ended concepts, eliminating the need for labeled data or human examples. This method integrates CLIP [27], which efficiently learns deep visual representations from natural language supervision. It utilizes both the image encoder and text encoder components of a CLIP model to compute the text embedding for each concept $s\in\mathcal{C}\subset\mathcal{S}$ from a concept dataset and the image embeddings for the probing images in the dataset, subsequently calculating a concept-activation matrix.
The activations of a target neuron $f_{i}$ are then computed across all images in the probing dataset $\boldsymbol{X}$. However, as this process is designed for scalar neurons, these activations are summarized by a function that calculates the mean of the activation map over spatial dimensions. The concept corresponding to the target neuron is determined by identifying the most similar concept $s$ based on its activation vector. The most highly activated images are denoted as $\boldsymbol{X}_{s}\subset\boldsymbol{X}$.
@@ A.1.6 INVERT
Labeling Neural Representations with Inverse Recognition (INVERT) [16] shares the capability of constructing complex explanations like CompExp [12] but with the added advantage of not relying on segmentation masks and only needing labeled data. The method obtains its explanations by merging individual concepts into logical formulas using composition operators AND, OR, and NOT. It also exhibits greater versatility in handling various neuron types and is computationally less demanding compared to previous methods such as NetDissect [11] and CompExp [12]. Additionally, INVERT introduces a transparent metric for assessing the alignment between representations and their associated explanations. The non-parametric Area Under the Receiver Operating Characteristic (AUC) measure evaluates the relationship between representations and concepts based on the representation’s ability to distinguish the presence from the absence of a concept, with statistical significance. The probing dataset with the concept present is labeled as $\boldsymbol{X}_{1}$, while the dataset without the concept is labeled as $\boldsymbol{X}_{0}$.
The goal of INVERT is to identify the concept $s\in\mathcal{L}\subset{S}$ that maximizes $\delta_{\text{AUC}}$ with the neuron $f_{i}$. Here, $s$ can be a combination of concepts. The optimization process resembles that of CompExp, employing beam search [49] to find the optimal compositional concept. The top-performing concepts are iteratively selected until the predefined compositional length $L\in\mathbb{N}$ is reached.
@@
The objective of INVERT is to maximize the similarity $\delta_{\text{AUC}}$ between a concept $s$ and the neuron $f_{i}$, which can be described as:
@@ A.2 Schematic Illustration of CoSy Implementation Details
In the example shown in Figure 1, we used the default settings of the explanation methods to generate explanations for neuron 80 in the avgpool layer of ResNet18. For CLIP-Dissect, we used the 20,000 most common English words as the concept dataset and the ImageNet validation dataset [36] as the probing dataset. We employed Stable Diffusion XL 1.0-base (SDXL[34] as the text-to-image model, using the prompt “realistic photo of a close up of [concept]” to generate concept images, with [concept] being replaced by the textual explanation from the methods. We generated 50 images per concept for 50 randomly chosen neurons from the avgpool layer of ResNet18. For evaluation, we also used the ImageNet validation dataset as the control dataset.
A.3 Prompt Bias Analysis
To address prompt bias and dataset dependency, we extended our analysis by comparing results on the object-focused ImageNet with the scene-focused Places365 dataset. Our results show a significant difference based on prompt selection and dataset: close-ups work well for object-centric datasets like ImageNet, while more general prompts, such as “photo of [concept]”, are more suitable for scene-based datasets like Places356 (see Figure 7).
Figure 7: Cosine Similarity between synthetic and natural concept images for 50 concepts in the scene-based dataset Places365 [42], using different input prompts for the text-to-image models Stable Diffusion XL 1.0-base (SDXL) and Stable Cascade. The third prompt, “photo of [concept]”, performs best for generating scene images.
Additionally, we evaluated model performance using different similarity metrics across both datasets. Alongside Cosine Similarity (CS), we introduced two more distance measures: Learned Perceptual Image Patch Similarity (LPIPS) [53], which calculates perceptual similarity between two images and aligns well with human perception, using deep embeddings from a VGG model. Additionally, we included “Euclidean Distance” (ED) to capture the absolute differences in pixel values. The results of these evaluations are presented in Table 5.
Table 5: Synthetic-to-Natural Image Similarity. This table illustrates the impact of varying parameters within our method pipeline. We evaluate five different prompts as input to two different text-to-image models, Stable Diffusion XL 1.0-base (SDXL) and Stable Cascade (SC). We randomly selected 50 from both ImageNet and Places365, using each class name as input to the prompt, denoted as [concept], generating 50 images per prompt for each model. We compute the average similarity metrics across all classes and report the standard deviation. Higher values of CS (Cosine Similarity) and lower values of ED (“Euclidean Distance”) and LPIPS (Learned Perceptual Image Patch Similarity) indicate greater similarity between synthetic and natural images. High similarity is desirable; bold values represent the highest scores for each prompt.
Dataset Prompt Text-to-image Similarity Measure
CS ($\uparrow$) ED ($\downarrow$) LPIPS ($\downarrow$)
ImageNet 1. “a [concept] SDXL 0.70$\pm$0.08 8.0$\pm$1.09 0.72$\pm$0.03
SC 0.69$\pm$0.08 8.19$\pm$1.02 0.74$\pm$0.04
2. “a painting of [concept] SDXL 0.64$\pm$0.07 8.74$\pm$0.81 0.73$\pm$0.03
SC 0.64$\pm$0.07 8.81$\pm$0.76 0.76$\pm$0.04
3. “photo of [concept] SDXL 0.71$\pm$0.08 7.94$\pm$1.11 0.72$\pm$0.03
SC 0.69$\pm$0.08 8.25$\pm$1.01 0.75$\pm$0.04
4. “realistic photo of [concept] SDXL 0.70$\pm$0.08 8.07$\pm$1.09 0.72$\pm$0.03
SC 0.69$\pm$0.08 8.34$\pm$0.99 0.73$\pm$0.03
5. “realistic photo of a close up of [concept] SDXL 0.72$\pm$0.08 7.97$\pm$1.11 0.72$\pm$0.03
SC 0.69$\pm$0.08 8.31$\pm$1.01 0.73$\pm$0.04
Places365 1. “a [concept] SDXL 0.62$\pm$0.09 9.02$\pm$1.04 0.71$\pm$0.02
SC 0.63$\pm$0.09 8.96$\pm$1.04 0.72$\pm$0.03
2. “a painting of [concept] SDXL 0.57$\pm$0.08 9.51$\pm$0.83 0.72$\pm$0.02
SC 0.58$\pm$0.08 9.43$\pm$0.83 0.72$\pm$0.03
3. “photo of [concept] SDXL 0.64$\pm$0.09 8.85$\pm$1.09 0.71$\pm$0.02
SC 0.64$\pm$0.09 8.89$\pm$1.08 0.72$\pm$0.03
4. “realistic photo of [concept] SDXL 0.62$\pm$0.09 9.08$\pm$1.02 0.71$\pm$0.02
SC 0.62$\pm$0.08 9.21$\pm$0.98 0.72$\pm$0.03
5. “realistic photo of a close up of [concept] SDXL 0.63$\pm$0.09 9.04$\pm$1.05 0.71$\pm$0.03
SC 0.63$\pm$0.09 9.14$\pm$1.01 0.72$\pm$0.03
A.4 Sanity Check Class Exclusion
In addition to the results presented in Table 2, we conducted the same experiment with the ground truth images excluded from the control dataset. The corresponding results, shown in Table 6, closely align with those obtained when the ground truth class was included. The findings are largely consistent with those obtained when the ground truth class was included.
Table 6: Comparison of true and random explanations on output neurons with known ground truth labels. This table presents the average quality scores (with standard deviation) for true explanations, derived from target class labels, and random explanations, derived from randomly selected synthetic image classes (excluding the target class), across four models pre-trained on ImageNet. Higher values are better. Our results consistently show high scores for true explanations and low scores for random ones.
Model AUC ($\uparrow$) MAD ($\uparrow$)
True Random True Random
ResNet18 0.98$\pm$0.09 0.52$\pm$0.24 6.59$\pm$2.14 0.08$\pm$0.91
DenseNet161 0.99$\pm$0.08 0.52$\pm$0.24 7.33$\pm$1.91 0.04$\pm$0.84
GoogLeNet 0.99$\pm$0.07 0.49$\pm$0.24 8.01$\pm$2.28 -0.01$\pm$0.81
ViT-B/16 0.99$\pm$0.05 0.52$\pm$0.22 14.7$\pm$3.88 0.13$\pm$1.11
@@ A.5 Intraclass Image Similarity
Cosine Similarity (CS), “Euclidean distance” (ED), and Learned Perceptual Image Patch Similarity (LPIPS) are commonly used metrics for measuring image similarity because they capture different aspects of similarity and complement each other. We compute the average CS, ED, and LPIPS for each class and determine the overall class average. Table 7 provides a detailed overview of the results quantifying the similarity within synthetic images using CS and ED. When evaluating these results, it is important to note that high scores do not necessarily indicate optimal outcomes, as they suggest nearly identical images, which may lack intraclass distance. Conversely, very low scores imply significant differences among images, which might not capture the essence of the concept adequately. Ideally, we aim for somewhat similar yet slightly varied images representing the same class. The results show that the Stable Cascade (SC) model consistently achieves higher scores across all prompts compared to the Stable Diffusion XL 1.0-base (SDXL) model. Notably, it obtains the highest score for the two most elaborate prompts (4, 5). This indicates that the SC model tends to offer less intraclass distance in visually representing concepts.
Table 7: Intraclass Image Similarity. This table illustrates the impact of varying parameters within our CoSy framework. We evaluate five different prompts as input to two different text-to-image models. A random selection of 50 ImageNet classes is made, with each class name used as input to the prompt, denoted as [concept], resulting in 50 images generated per prompt using a text-to-image model. We compute the average intraclass similarity across all classes and report the standard deviation. Higher CS and lower ED values indicate greater similarity between the images. In intraclass image similarity, neither excessively high nor excessively low scores are desirable.
Dataset Prompt Text-to-image Similarity Measure
CS ($\uparrow$) ED ($\downarrow$) LPIPS ($\downarrow$)
ImageNet 1. “a [concept] SDXL 0.85$\pm$0.07 5.61$\pm$1.40 0.61$\pm$0.11
SC 0.93$\pm$0.03 3.76$\pm$0.92 0.49$\pm$0.10
2. “a painting of [concept] SDXL 0.87$\pm$0.05 4.94$\pm$1.13 0.62$\pm$0.10
SC 0.93$\pm$0.03 3.67$\pm$0.87 0.54$\pm$0.10
3. “photo of [concept] SDXL 0.84$\pm$0.06 5.65$\pm$1.34 0.62$\pm$0.10
SC 0.92$\pm$0.04 4.03$\pm$0.99 0.52$\pm$0.10
4. “realistic photo of [concept] SDXL 0.87$\pm$0.06 5.15$\pm$1.31 0.60$\pm$0.10
SC 0.94$\pm$0.03 3.53$\pm$0.86 0.45$\pm$0.09
5. “realistic photo of a close up of [concept] SDXL 0.89$\pm$0.04 4.82$\pm$1.16 0.60$\pm$0.10
SC 0.94$\pm$0.03 3.58$\pm$0.86 0.50$\pm$0.09
Places365 1. “a [concept] SDXL 0.84$\pm$0.06 5.75$\pm$1.35 0.61$\pm$0.09
SC 0.92$\pm$0.03 4.02$\pm$0.95 0.54$\pm$0.09
2. “a painting of [concept] SDXL 0.88$\pm$0.04 4.89$\pm$1.05 0.61$\pm$0.09
SC 0.93$\pm$0.03 3.79$\pm$0.84 0.56$\pm$0.09
3. “photo of [concept] SDXL 0.85$\pm$0.06 5.61$\pm$1.27 0.61$\pm$0.10
SC 0.93$\pm$0.03 3.94$\pm$0.92 0.54$\pm$0.09
4. “realistic photo of [concept] SDXL 0.88$\pm$0.05 5.05$\pm$1.15 0.59$\pm$0.09
SC 0.94$\pm$0.03 3.70$\pm$0.88 0.52$\pm$0.08
5. “realistic photo of a close up of [concept] SDXL 0.87$\pm$0.05 5.25$\pm$1.24 0.60$\pm$0.09
SC 0.94$\pm$0.03 3.68$\pm$0.91 0.54$\pm$0.09